Tickets for a basketball tournament were $6 for students and $9 for nonstudents. Total sales were $10,500, and 250 more student tidkets were sold than nonstudents. How many of each type of ticket were sold?
Answer by texttutoring(324)

There are two variables:
Let x = the number of $6 tickets
Let y = the number of $9 tickets

You need two equations.

Total sales were $10,500, so

Equation 1: 6x +9y = 10500

There were 250 more student tickets sold than nonstudent tickets, so

Equation 2: x = y + 250

Now take x=y+250 from Equation 2 and substitute it into Equation 1:

6(y+250) +9y = 10500
6y+1500 +9y = 10500
15y = 10500-1500
15y=9000
y=600

Now find x:

x=y+250
x=600+250
x=850

There were 850 student tickets and 600 nonstudent tickets.